Which of the following is a linear equation in x; y and z? 1. x 2 − e 2y = 3z, where e = 2.71828 .... 2. 2π ln(e z ) − 2y + z = ln(3) − x. 3. p y −1 + 4y − 2z = 7x. 4. y + 4y − 2z = 7x 3 .


1
Expert's answer
2021-05-17T01:56:02-0400

The equation "x-e^2 y=3z" is a linear equation in x, y and z because-

      

   Since, "x-e^2y-3z=0"


here, "a=1,b=-e^2,c=-3"

      Here "a\neq b\neq c\neq 0"

 

   SO It is a linear equation in x, y and z.